OpenStudy (azionne):

7(x -2) + 3(x + 2) = 5(x - 6) will medal

pooja195 (pooja195):

Start by distributing. \[\rm~(7)(x)+(7)(−2)+(3)(x)+(3)(2)=(5)(x)+(5)(−6)\] Write the equation you get.

OpenStudy (azionne):

@bbypanda16

pooja195 (pooja195):

@Azionne you are simply trying to find answers. I would appreciate if you did the work rather than look for answers.

OpenStudy (azionne):

i aint telling them to give me answers so id appreciate it if u would back off

pooja195 (pooja195):

They gave you multiple answers in the past. It's why u tagged them.

OpenStudy (igigighjkl):

?

OpenStudy (azionne):

not tru they r the best help and i dont think ur good help

pooja195 (pooja195):

Yeah the difference is I try to teach you they simply explain and give the answer.

OpenStudy (azionne):

its the way i understand thank u very much

OpenStudy (igigighjkl):

Uh... me?

OpenStudy (azionne):

no im talking to pooja195

pooja195 (pooja195):

Ok. Have fun. I'll let them help :) Bye have a nice day.

OpenStudy (azionne):

yea u to

OpenStudy (anonymous):

u wont get long this way buddy :) remember these words :) All the best!!!

OpenStudy (azionne):

^ what

pooja195 (pooja195):

We would have been done with the question by now.

TheSmartOne (thesmartone):

@tkhunny we found someone who doesn't want to use openstudy for what it is meant for :)

OpenStudy (azionne):

lol whos that @TheSmartOne

OpenStudy (anonymous):

you will understand this better: solve 3(x -2) + 7(x + 2) = 4(x - 6)+(x-6)

OpenStudy (azionne):

thx^

OpenStudy (anonymous):

Solve the above question since you learned it a while ago in your last problem how to solve....

TheSmartOne (thesmartone):

if you solve that, you're going to get a wrong answer

TheSmartOne (thesmartone):

the actual equation is 7(x -2) + 3(x + 2) = 5(x - 6

pooja195 (pooja195):

Start by distributing. \[\rm~(7)(x)+(7)(−2)+(3)(x)+(3)(2)=(5)(x)+(5)(−6)\] Write the equation you get.

OpenStudy (anonymous):

@TheSmartOne -.-

TheSmartOne (thesmartone):

The question states 3(x+2) and you gave them 3(x-2) You're definitely going to get a wrong answer. Same, question says 7(x-2), but you changed it to 7(x+2)

OpenStudy (anonymous):

i changed the question for him to answer since he declares that he learns it way better than how @pooja195 was teaching him..

TheSmartOne (thesmartone):

but you changed subtraction to addition, and vice versa and that will change the solution too :)

OpenStudy (anonymous):

yea so what if he learnt how to solve in the last problem he must be able to solve this question too...

OpenStudy (anonymous):

anyways i leave it to the mods..

pooja195 (pooja195):

He doesn't like the way I taught because I don't give answers.

OpenStudy (anonymous):

@Azionne if u learnt how to tackle such type of question then solve this right now : 5(x+2)-7(x+5)=2

OpenStudy (anonymous):

i hope u learn what i mean by this...

OpenStudy (azionne):

lol 5x+10-7-35=2

OpenStudy (anonymous):

sure?

OpenStudy (pawanyadav):

If I correct him I will get another warning

OpenStudy (tkhunny):

You're missing an 'x'.

OpenStudy (anonymous):

i want the user to understand that direct answers wont lead him anywhere other than these marks.. Marks will lead u to job bascially wherein u have to apply you're own mind instead of receiving the final answer try to understand it maybe you will learn some new step to tackle the same problem